# Second-Order Fine-Tuning without Pain for LLMs: A Hessian Informed Zeroth-Order Optimizer

### v1

@@ Appendix B Detailed Convergence Analysis
$=$ $\mathbb{E}\bigg{[}\|\frac{1}{b}\sum_{i=1}^{b}\Sigma_{t}^{1/2}u_{i}u_{i}^{\top}\Sigma_{t}^{1/2}\nabla\mathcal{L}(\theta_{t},\mathcal{B}_{t})+\mathcal{O}(\mu)\|^{2}\bigg{]}$
$\leq$ $2\mathbb{E}\left[\|\frac{1}{b}\sum_{i=1}^{b}\Sigma_{t}^{1/2}u_{i}u_{i}^{\top}\Sigma_{t}^{1/2}\nabla\mathcal{L}(\theta_{t},\mathcal{B}_{t})\|^{2}\right]+\mathcal{O}(\mu^{2})$
$\leq$ $\frac{2}{b}\sum_{i=1}^{b}\mathbb{E}\left[\|\Sigma_{t}^{1/2}u_{i}u_{i}^{\top}\Sigma_{t}^{1/2}\nabla\mathcal{L}(\theta_{t},\mathcal{B}_{t})\|^{2}\right]+\mathcal{O}(\mu^{2})$
$=$ $2\mathrm{tr}(\Sigma_{t})\cdot\nabla^{\top}\mathcal{L}(\theta_{t},\mathcal{B}_{t})\Sigma_{t}\nabla\mathcal{L}(\theta_{t},\mathcal{B}_{t})$
$+2\nabla^{\top}\mathcal{L}(\theta_{t},\mathcal{B}_{t})\Sigma_{t}^{2}\nabla\mathcal{L}(\theta_{t},\mathcal{B}_{t})+\mathcal{O}(\mu^{2})$
$\leq$ $2\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\nabla^{\top}\mathcal{L}(\theta_{t},\mathcal{B}_{t})\Sigma_{t}\nabla\mathcal{L}(\theta_{t},\mathcal{B}_{t})+\mathcal{O}(\mu^{2}),$
@@
$\mathbb{E}\left[\nabla^{\top}\mathcal{L}(\theta_{t},\mathcal{B}_{t})\Sigma_{t}\nabla\mathcal{L}(\theta_{t},\mathcal{B}_{t})\right]=\mathbb{E}\left[\|\nabla\mathcal{L}(\theta_{t},\mathcal{B}_{t})\|_{\Sigma_{t}}^{2}\right]$
$\leq$ $2\mathbb{E}\left[\|\nabla\mathcal{L}(\theta_{t})\|_{\Sigma_{t}}^{2}\right]+2\mathbb{E}\left[\|\nabla\mathcal{L}(\theta_{t},\mathcal{B}_{t})-\nabla\mathcal{L}(\theta_{t})\|_{\Sigma_{t}}^{2}\right]$
$\leq$ $2\|\nabla\mathcal{L}(\theta_{t})\|_{\Sigma_{t}}^{2}+2\beta_{u}\mathbb{E}\left[\|\nabla\mathcal{L}(\theta_{t},\mathcal{B}_{t})-\nabla\mathcal{L}(\theta_{t})\|^{2}\right]$
@@
where the second inequality is because of Assumption 4.3 and the last inequality is because of Assumption 4.2.
@@
(Magnus et al., 1978) Let $A$ and $B$ be two symmetric matrices, and $u$ obeys the Gaussian distribution, that is, $u\sim N(0,I_{d})$. Define $z=u^{\top}Au\cdot u^{\top}Bu$. The expectation of $z$ is:
@@
Appendix C MeZO Variants
There is a rich history of transferring ideas from first order optimization to enhance ZO algorithms. Below, we highlight the variant of HiZOO: $n-$HiZOO which can perform $n$ estimation times per step efficiently as shown in Algorithm 2. In particular, if sampling $n$ $u$ vectors and averaging the projected gradients, it can decrease the variance of the diagonal Hessian matrix and reduce the overall training steps, but will cause much more computation per step meanwhile.
@@ Appendix C MeZO Variants
  Input: parameters $\theta\in\mathbb{R}^{d}$, loss $L:\mathbb{R}^{d}\rightarrow\mathbb{R}$, step budget $T$, perturbation scale $\mu$, batch size $B$, learning rate schedule ${\eta_{t}}$, smooth scale $\alpha_{t}$, Hessian matrix $\Sigma_{0}$
  for $t=1,...,T$ do
     seeds, projected_grads $\leftarrow$ []
     for $j=1,...,b$ do
        Sample batch $\mathcal{B}\subset\mathcal{D}$ and random seed $s$
        $\ell\leftarrow\mathcal{L}(\theta;\mathcal{B})$
        $\theta$leftarrow PerturbParameters($\theta$, $\mu$, $\Sigma^{1/2}_{t-1}$, $s$)
        $\ell_{+}\leftarrow\mathcal{L}(\theta;\mathcal{B})$
        $\theta$leftarrow PerturbParameters($\theta$, $-2\mu$, $\Sigma^{1/2}_{t-1}$, $s$)
        $\ell_{-}\leftarrow\mathcal{L}(\theta;\mathcal{B})$
@@
        $\theta$leftarrow PerturbParameters($\theta$, $\mu$, $\Sigma^{1/2}_{t-1}$, $s$) //Reset parameters before descent
        $\Sigma^{\prime}_{t}=\frac{1}{2\mu^{2}}(\ell_{+}+\ell_{+}-2\ell)(\Sigma^{-1/2}_{t-1}u_{i}u_{i}^{\top}\Sigma^{-1/2}_{t-1})$
@@
        $\Sigma_{t}=(1-\alpha_{t})\Sigma_{t-1}+\alpha_{t}\left|diag(\Sigma^{\prime}_{t})\right|$ //Update Hessian matrix
        projected_grad $\leftarrow$ $(\ell_{+}-\ell_{-})/2\mu$ projected_grads[j] $\leftarrow$ projected_grad seeds[j] $\leftarrow s$
     end for
  end for
  for $j=1,...,b$ do
     Reset random number generator with seeds[j]
     for $\theta_{i}\in\theta$ do
        $z\sim\mathcal{N}(0,I_{d})$ $*\theta_{i}\leftarrow\theta_{i}-(\eta_{t}/n)*$projected_grads[j] *$*z$ //Avg grad for $u_{1},...,u_{n}$
     end for
  end forPerturbParameters($\theta$, $\mu$, $\Sigma^{1/2}_{t}$,$s$):
@@
  Reset random number generator with seed $s$
  for $\theta_{i}\in\theta$  do
     $u_{i}\sim\mathcal{N}(0,I_{d})$
     $\theta_{i}\leftarrow\theta_{i}+\mu\Sigma^{1/2}_{t}u_{i}$ Modify parameters in place
     Return $\theta$
  end for
Algorithm 2 HiZOO with $b>1$
Appendix D Details about Experiments on RoBERTa-large
D.1 Experiment Hyperparameters
@@
D.2 Experiment Results
Here we show the full experiment results in Table 7 and plot more loss curves to compare with MeZO. As shown in Figure 9 and Figure 8, we can see that HiZOO can greatly accelerate the training process over MeZO, which verifies the robustness of HiZOO.
Task Type SST-2 SST-5 SNLI MNLI RTE TREC
—— sentiment —— —— natural language inference —— — topic —
Zero-shot 79.0 35.5 50.2 48.8 51.4 32.0
LP 91.3 ($\pm$0.5) 51.7 ($\pm$0.5) 80.9 ($\pm$1.0) 71.5 ($\pm$1.1) 73.1 ($\pm$1.5) 89.4 ($\pm$0.5)
FT 91.9 ($\pm$1.8) 47.5 ($\pm$1.9) 77.5 ($\pm$2.6) 70.0 ($\pm$2.3) 66.4 ($\pm$7.2) 85.0 ($\pm$2.5)
FT (LoRA) 91.4 ($\pm$1.7) 46.7 ($\pm$1.1) 74.9 ($\pm$4.3) 67.7 ($\pm$1.4) 66.1 ($\pm$3.5) 82.7 ($\pm$4.1)
FT (prefix) 91.9 ($\pm$1.0) 47.7 ($\pm$1.1) 77.2 ($\pm$1.3) 66.5 ($\pm$2.5) 66.6 ($\pm$2.0) 85.7 ($\pm$1.3)
MeZO 93.3 ($\pm$0.7) 53.2 ($\pm$1.4) 83.0 ($\pm$1.0) 78.3 ($\pm$0.5) 78.6 ($\pm$2.0) 94.3 ($\pm$1.3)
MeZO (LoRA) 93.4 ($\pm$0.4) 52.4 ($\pm$0.8) 84.0 ($\pm$0.8) 77.9 ($\pm$0.6) 77.6 ($\pm$1.3) 95.0 ($\pm$0.7)
MeZO (prefix) 93.3 ($\pm$0.1) 53.6 ($\pm$0.5) 84.8 ($\pm$1.1) 79.8 ($\pm$1.2) 77.2 ($\pm$0.8) 94.4 ($\pm$0.7)
HiZOO 95.5 ($\pm$0.4) 52.8 ($\pm$0.9) 82.6 ($\pm$0.7) 76.0 ($\pm$0.6) 80.0 ($\pm$ 1.5) 94.6 ($\pm$1.1)
HiZOO(LoRA) 91.7 ($\pm$0.3) 45.3 ($\pm$0.7) 76.5 ($\pm$0.3) 63.1 ($\pm$0.6) 70.4 ($\pm$1.4) 59.6 ($\pm$1.5)
HiZOO(prefix) 96.1 ($\pm$0.2) 54.2 ($\pm$0.4) 85.7 ($\pm$0.7) 79.7 ($\pm$1.0) 77.3 ($\pm$0.2) 93.9 ($\pm$0.6)
Table 7: Experiments on RoBERTa-large (350M parameters, k=512) that include zero-shot learning, linear probing (LP), full-parameter fine-tuning with Adam, MeZO and HiZOO, and parameter-efficient Fine-tuning (LoRA and prefix learning) with Adam, MeZO and HiZOO respectively. All reported numbers are averaged accuracy (standard deviation) over 5 runs.
Figure 8: Experiments on RTE with 3 fine-tuning methods.
Figure 9: Experiments on SNLI with 3 fine-tuning methods.
Appendix E Details about Experiments on OPT
We use the hyperparameters in Table 8 for HiZOO experiments on OPT. We also provide more loss curves of fine-tuning OPT family in Figure 10.
@@ Appendix E Details about Experiments on OPT
Figure 10: More experiment results on OPT.
@@ Appendix F Details about Memory Usage
Here we show the detailed numbers of memory profiling results Table 9, which also corresponds to Table 5. We did not turn on any advance memory-saving options, e.g., gradient checkpointing. We set the per-device batch size as 1 to test the minimum hardware requirement to run the model with specific optimization algorithms. We use Nvidia’s $nvidia-smi$ command to monitor the GPU memory usage.
Method MeZO HiZOO ICL Prefix FT Full-parameter FT
1.3B 1xA100 (4GB) 1xA100 (7GB) 1xA100 (6GB) 1xA100 (19GB) 1xA100 (27GB)
2.7B 1xA100 (7GB) 1xA100 (13GB) 1xA100 (8GB) 1xA100 (29GB) 1xA100 (55GB)
6.7B 1xA100 (14GB) 1xA100 (29GB) 1xA100 (16GB) 1xA100 (46GB) 2xA100 (156GB)
13B 1xA100 (26GB) 1xA100 (53GB) 1xA100 (29GB) 2xA100 (158GB) 4xA100 (316GB)
30B 1xA100 (58GB) 1xA100 (118GB) 1xA100 (62GB) 4xA100 (315GB) 8xA100 (633GB)
66B 2xA100 (128GB) 1xA100 (258GB) 2xA100 (134GB) 8xA100 16xA100
Table 9: Memory usage on the MultiRC (average tokens=400) dataset. Results of ICL,Prefix FT and Full-parameter FT are from MeZO(Malladi et al., 2023).
Appendix G Details about Ablation Experiments
G.1 Influence of Estimation Times $n$ Per Step
We conducted experiments to explore the influence of estimation times $n$ per step as shown in Figure 11. We can conclude that when $n$ is larger, the estimation of diagonal Hessian is more accurate, finding the direction of gradient descent more precisely. But it requires more computation with $n$ times per step. So choosing an appropriate value of $n$ is very important during the training.
Figure 11: Influence of estimation times per step. (a) 2D trajectories of gradient descent; (b) Corresponding loss curves.
G.2 Influence of Smooth Scale $\alpha_{t}$
We conducted experiments on SST-2, SST-5, MNLI datasets when fine-tuning RoBERTa-large to research the influence of smooth scale $\alpha_{t}$. Figure 5 shows that the value of $n$ mainly affects the convergence speed of the model, but has little effect on the final result of the training. Additionally, the best value of $n$ will vary between different datasets.

### v3

@@ Appendix B Detailed Convergence Analysis
$=$ $\mathbb{E}\bigg{[}\|\frac{1}{b}\sum_{i=1}^{b}\Sigma_{t}^{1/2}u_{i}u_{i}^{\top}\Sigma_{t}^{1/2}\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})+\mathcal{O}(\mu)\|^{2}\bigg{]}$
$\leq$ $2\mathbb{E}\left[\|\frac{1}{b}\sum_{i=1}^{b}\Sigma_{t}^{1/2}u_{i}u_{i}^{\top}\Sigma_{t}^{1/2}\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})\|^{2}\right]+\mathcal{O}(\mu^{2})$
$\leq$ $\frac{2}{b}\sum_{i=1}^{b}\mathbb{E}\left[\|\Sigma_{t}^{1/2}u_{i}u_{i}^{\top}\Sigma_{t}^{1/2}\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})\|^{2}\right]+\mathcal{O}(\mu^{2})$
$=$ $2\mathrm{tr}(\Sigma_{t})\cdot\nabla^{\top}\mathcal{L}(\theta_{t};\mathcal{B}_{t})\Sigma_{t}\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})$
$+2\nabla^{\top}\mathcal{L}(\theta_{t};\mathcal{B}_{t})\Sigma_{t}^{2}\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})+\mathcal{O}(\mu^{2})$
$\leq$ $2\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\nabla^{\top}\mathcal{L}(\theta_{t};\mathcal{B}_{t})\Sigma_{t}\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})+\mathcal{O}(\mu^{2}),$
@@
$\mathbb{E}\left[\nabla^{\top}\mathcal{L}(\theta_{t};\mathcal{B}_{t})\Sigma_{t}\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})\right]=\mathbb{E}\left[\|\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})\|_{\Sigma_{t}}^{2}\right]$
$\leq$ $2\mathbb{E}\left[\|\nabla\mathcal{L}(\theta_{t})\|_{\Sigma_{t}}^{2}\right]+2\mathbb{E}\left[\|\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})-\nabla\mathcal{L}(\theta_{t})\|_{\Sigma_{t}}^{2}\right]$
$\leq$ $2\|\nabla\mathcal{L}(\theta_{t})\|_{\Sigma_{t}}^{2}+2\beta_{u}\mathbb{E}\left[\|\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})-\nabla\mathcal{L}(\theta_{t})\|^{2}\right]$
@@
where the second inequality is because of Assumption B.3 and the last inequality is because of Assumption B.2.
@@
[30] Let $A$ and $B$ be two symmetric matrices, and $u$ obeys the Gaussian distribution, that is, $u\sim N(0,I_{d})$. Define $z=u^{\top}Au\cdot u^{\top}Bu$. The expectation of $z$ is:
@@
Appendix C Test Functions of the optimization trajectories
For better illustrating how HiZOO utilizes hessian to improve the convergence process, we choose below three test functions with heterogeneous curvatures across different parameters. In Figure 8, we provide the 2D convergence paths of three functions and the variation of their losses with respect to steps.
  • Function (a) [^1] Function (a) is from [26].: $f(x,y)=8(x-1)^{2}(1.3x^{2}+2x+1)+0.5(y-4)^{2}$
  • Function (b): $f(x,y)=\lvert x\rvert+\lvert y\rvert$
  • Function (c): $f(x,y)=10000x^{2}+y^{2}$
Figure 9: Influence of estimation times per step. (left) 2D trajectories of gradient descent; (right) Corresponding loss curves.
@@ Appendix C Test Functions of the optimization trajectories
1:parameters $\theta\in\mathbb{R}^{d}$, loss $L:\mathbb{R}^{d}\rightarrow\mathbb{R}$, step budget $T$, perturbation scale $\mu$, batch size $B$, learning rate schedule ${\eta_{t}}$, smooth scale $\alpha_{t}$, estimate times $n$, Hessian matrix $\Sigma_{0}$
2:for $t=1,...,T$ do
3:     seeds, projected_grads $\leftarrow$ []
4:     for $j=1,...,n$ do
5:         Sample batch $\mathcal{B}\subset\mathcal{D}$ and random seed $s$
6:         $\ell\leftarrow\mathcal{L}(\theta;\mathcal{B})$
7:         $\theta$leftarrow PerturbParameters($\theta$, $\mu$, $\Sigma^{1/2}_{t-1}$, $s$)
8:         $\ell_{+}\leftarrow\mathcal{L}(\theta;\mathcal{B})$
9:         $\theta$leftarrow PerturbParameters($\theta$, $-2\mu$, $\Sigma^{1/2}_{t-1}$, $s$)
10:         $\ell_{-}\leftarrow\mathcal{L}(\theta;\mathcal{B})$
@@
11:         $\theta$leftarrow PerturbParameters($\theta$, $\mu$, $\Sigma^{1/2}_{t-1}$, $s$) $\triangleright$ Reset parameters before descent
12:         $\Sigma^{\prime}_{t}=\frac{1}{2\mu^{2}}(\ell_{+}+\ell_{+}-2\ell)(\Sigma^{-1/2}_{t-1}u_{i}u_{i}^{\top}\Sigma^{-1/2}_{t-1})$
@@
13:         $\Sigma_{t}^{-1}=(1-\alpha_{t})\Sigma_{t-1}^{-1}+\alpha_{t}\left|diag(\Sigma^{\prime}_{t})\right|$triangleright Update Hessian matrix
14:         projected_grad $\leftarrow$ $(\ell_{+}-\ell_{-})*\Sigma^{1/2}_{t}/2\mu$
15:         projected_grads[j] $\leftarrow$ projected_grad
16:         seeds[j] $\leftarrow s$
17:     end for
18:end for
19:for $j=1,...,n$ do
20:     Reset random number generator with seeds[j]
21:     for $\theta_{i}\in\theta$ do
22:         $u_{i}\sim\mathcal{N}(0,I_{d})$
23:         $\theta_{i}\leftarrow\theta_{i}-(\eta_{t}/n)*$projected_grads[j] $*u_{i}$triangleright Avg grad for $u_{1},...,u_{n}$
24:     end for
25:end for
26:
27:function PerturbParameter($\theta$, $\mu$, $\Sigma^{1/2}_{t}$,$s$)
28:     Reset random number generator with seed $s$
29:     for $\theta_{i}\in\theta$  do
30:         $u_{i}\sim\mathcal{N}(0,I_{d})$
31:         $\theta_{i}\leftarrow\theta_{i}+\mu\Sigma^{1/2}_{t}u_{i}$triangleright Modify parameters in place
32:     end for
33:     return $\theta$
34:end function
Algorithm 2 HiZOO-multi
1:parameters $\theta\in\mathbb{R}^{d}$, loss $L:\mathbb{R}^{d}\rightarrow\mathbb{R}$, step budget $T$, perturbation scale $\mu$, learning rate schedule ${\eta_{t}}$, smooth scale $\alpha_{t}$, diagonal Hessian $R_{0}$, $C_{0}$
2:for $t=1,...,T$ do
3:     Sample batch $\mathcal{B}\subset\mathcal{D}$ and random seed $s$
4:     $\ell\leftarrow\mathcal{L}(\theta;\mathcal{B})$
5:     $\theta$leftarrow PerturbParameters($\theta$, $\mu$, $R_{t-1}$, $C_{t-1}$, $s$)
6:     $\ell_{+}\leftarrow\mathcal{L}(\theta;\mathcal{B})$
7:     $\theta$leftarrow PerturbParameters($\theta$, $-2\mu$, $R_{t-1}$, $C_{t-1}$, $s$)
8:     $\ell_{-}\leftarrow\mathcal{L}(\theta;\mathcal{B})$
9:     $\theta$leftarrow PerturbParameters($\theta$, $\mu$, $R_{t-1}$, $C_{t-1}$, $s$) $\triangleright$ Reset parameters before descent
10:     $\hat{\Sigma}^{-1}_{t-1}=(R_{t-1}*C_{t-1})/(1_{n}^{\top}*R_{t-1})$
11:     $\hat{\Sigma}^{\prime}_{t}=\frac{1}{2\mu^{2}}(\ell_{+}+\ell_{-}-2\ell)(\hat{\Sigma}^{-1/2}_{t-1}u_{i}u_{i}^{\top}\hat{\Sigma}^{-1/2}_{t-1})$
12:     $R_{t}^{-1}=(1-\alpha_{t})R_{t-1}^{-1}+\alpha_{t}\left|diag(\hat{\Sigma}^{\prime}_{t})\right|*1_{m}$
13:     $C_{t}^{-1}=(1-\alpha_{t})C_{t-1}^{-1}+\alpha_{t}1^{\top}_{n}*\left|diag(\hat{\Sigma}^{\prime}_{t})\right|$
14:     projected_grad $\leftarrow(\ell_{+}-\ell_{-})\ast\hat{\Sigma}^{1/2}_{t}/2\mu$
@@
15:     Reset random number generator with seed $s$triangleright For sampling $u_{i}$
16:     for $\theta_{i}\in\theta$ do
17:         Sample $u_{i}\sim\mathcal{N}(0,I_{d})$
18:         $\theta_{i}\leftarrow\theta_{i}-\eta_{t}\ast$ projected_grad $\ast u_{i}$
19:     end for
20:end for
21:
22:function PerturbParameter($\theta$, $\mu$, $R_{t}$, $C_{t}$, $s$)
23:     Reset random number generator with seed $s$triangleright For sampling $u_{i}$
24:     for $\theta_{i}\in\theta$ do
25:         Sample $u_{i}\sim\mathcal{N}(0,I_{d})$
26:         $\hat{\Sigma}^{-1}_{t}=(R_{t}*C_{t})/(1_{n}^{\top}*R_{t})$
27:         $\theta_{i}\leftarrow\theta_{i}+\mu\hat{\Sigma}^{1/2}_{t}u_{i}$triangleright Modify parameters in place
28:     end for
29:     return $\theta$
30:end function
Algorithm 3 HiZOO-L
Appendix D HiZOO Variants
D.1 HiZOO-multi
There is a rich history of transferring ideas from first order optimization to enhance ZO algorithms. Below, we highlight the variant of HiZOO: HiZOO-multi which can perform $n$ estimation times per step efficiently as shown in Algorithm 2. We conducted experiments to explore the influence of estimation times $n$ per step as shown in Figure 9. We can conclude that when $n$ is larger, the estimation of diagonal Hessian is more accurate. It can decrease the variance of the estimated diagonal Hessian matrix during each step and thus reduce the overall training steps, but will cause much more computation per step meanwhile. So choosing an appropriate value of $n$ is very important during the training.
D.2 HiZOO-L
Due to the storage of Hessian, HiZOO introduces extra memory cost, which is equal to the size of the model parameters. To address this limitations, we propose HiZOO-L, the low-rank implementation for the storage of Hessian, motivated by Adafactor [41]. Details can be see in Algorithm 2. We conduct lots of experiments to show the performance of HiZOO-L, as results shown in Figure 10.
Figure 10: Accuracy comparison on OPT-13B (left) and Llama3 (right) between HiZOO and HiZOO-L.
Appendix E Experiments on LLMs
E.1 Detailed Experiments on RoBERTa-large
@@
Figure 11: Loss curves on RoBERTa-large between MeZO and HiZOO.
Here we plot more loss curves to compare with MeZO. As shown in Figure 11, we can see that HiZOO can greatly accelerate the training process over MeZO, which verifies the robustness of HiZOO.
E.2 Detailed results on OPT
We use the hyperparameters in Table 7 for HiZOO experiments on OPT. We also provide the relative loss curves of fine-tuning OPT family in Figure 12.
@@ E.2 Detailed results on OPT
Figure 12: Loss curves on OPT between MeZO and HiZOO.
E.3 Detailed results on Phi-2 and Llama3
We provide several loss curves of fine-tuning Phi-2(2.7B) and Llama3(8B) in Figure 13 and Figure 14.
Figure 13: Loss curves on Phi-2 between MeZO and HiZOO.
Figure 14: Loss curves on Llama3 between MeZO and HiZOO.
@@ Appendix F Details about Memory Usage
Here we show the detailed numbers of memory profiling results Table 8. We did not turn on any advance memory-saving options, e.g., gradient checkpointing. We set the per-device batch size as 1 to test the minimum hardware requirement to run the model with specific optimization algorithms. We use Nvidia’s $nvidia-smi$ command to monitor the GPU memory usage.
Method zero-shot/MeZO(FT) HiZOO(FT) HiZOO-L(FT) ICL Adam(FT)
1.3B 1xA100 (4GB) 1xA100 (7GB) 1xA100 (4GB) 1xA100 (6GB) 1xA100 (27GB)
2.7B 1xA100 (7GB) 1xA100 (13GB) 1xA100 (8GB) 1xA100 (8GB) 1xA100 (55GB)
6.7B 1xA100 (14GB) 1xA100 (29GB) 1xA100 (15GB) 1xA100 (16GB) 2xA100 (156GB)
13B 1xA100 (26GB) 1xA100 (53GB) 1xA100 (29GB) 1xA100 (29GB) 4xA100 (316GB)
30B 1xA100 (58GB) 2xA100 (118GB) 1xA100 (64GB) 1xA100 (62GB) 8xA100 (633GB)
66B 2xA100 (128GB) 3xA100 (246GB) 2xA100 (140GB) 2xA100 (134GB) 16xA100
Table 8: Memory usage on the MultiRC (average tokens=400) dataset. Results of ICL and full-parameter tuning are from MeZO[32].
Appendix G Details about Wallclock Time Efficiency
In this section, we measure the wallclock time efficiency of HiZOO compared to MeZO and full-parameter fine-tuning (FT) with respect to different model sizes. Due to the lack of NV-Link connectivity in our A100 GPUs, we selected models that can be fully fine-tuned on a single A100 GPU for comparison. As shown in Table 9, HiZOO exhibits a longer per-step duration compared to MeZO, within a 50% margin. This result indicates that the primary overhead in hierarchical optimization methods lies in the forward propagation process. Given that HiZOO involves an additional forward pass compared to MeZO, the time per step increases by approximately 1.4 to 1.5 times.
In conclusion, the speedup factors derived from the forward pass step used in our comparisons between HiZOO and MeZO reflect the actual wallclock time efficiency improvements accurately.
Model RoBERTa-large(350M) Phi-2(2.7B) Llama3(8B) OPT(13B)
MeZO 0.2092s(BS=64) 0.3011s(BS=16) 0.7471s(BS=16) 1.1108s(BS=16)
HiZOO 0.3023s(BS=64) 0.4486s(BS=16) 1.1090s(BS=16) 1.5225s(BS=16)
HiZOO-L 0.3193s(BS=64) 0.4851s(BS=16) 1.1996s(BS=16) 1.6422s(BS=16)
Table 9: Wallclock time per step between MeZO and HiZOO. The increase in wallclock time per step for HiZOO compared to MeZO is less than 1.5 times across different model sizes. To avoid introducing additional overheads such as inter-GPU communication, results are measured on the same dataset (SST-2) and GPUs (80GB A100), with each result averaged over 100 steps. "BS" refers to batch size. For the relatively smaller RoBERTa-large model, we used a BS=64, while for models larger than 1B parameters, we used a BS=16.
Appendix H Details about Ablation Experiments
H.1 Influence of Smooth Scale $\alpha_{t}$
We conducted experiments on SST-2, SST-5, MNLI datasets when fine-tuning RoBERTa-large to research the influence of smooth scale $\alpha_{t}$. Figure 15 shows that the value of $\alpha_{t}$ mainly affects the convergence speed of the model. Additionally, the best value of $\alpha_{t}$ will vary between different datasets.